Currently, the dividend-payout ratio (D/E) for the aggregate market is 60 percent, the required return (k) is 6 percent, and the expected growth rate for dividends (g) is 2 percent. Round your answers to two decimal places.

  1. Compute the current earnings multiplier.

    x

  2. You expect the D/E payout ratio to decline to 45 percent, but you assume there will be no other changes. What will be the P/E?

    x
Answer #1

a). Earnings Multiplier = D/E / [k - g]

= 0.6 / [0.06 - 0.02] = 0.6 / 0.04 = 15 x

b). P/E = D/E / [k - g]

= 0.45 / [0.06 - 0.02] = 0.45 / 0.04 = 11.25 x